OpenStudy (anonymous):

find the max/min/saddle points f(x,y) = x/y + 1/x + 1/y

OpenStudy (anonymous):

same thing? how do i find critical points?

OpenStudy (zarkon):

did you find fx and fy?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

fx=y - 1/x^2 fy=x - 1/y^2

OpenStudy (anonymous):

but when i set them equal, it doesn't give me anything

OpenStudy (zarkon):

i don't like your fy

OpenStudy (zarkon):

or your fx

OpenStudy (anonymous):

wait wait sorry the original equation is xy + 1/x + 1/y not x/y

OpenStudy (zarkon):

ok

OpenStudy (anonymous):

so if i set them equal, x=y and then x and y can be -1 and 0?

OpenStudy (zarkon):

\[y-\frac{1}{x^2}=0\] \[y=\frac{1}{x^2}\] and so \[x-\frac{1}{y^2}=0\] \[x-\frac{1}{1/x^4}=0\] \[x-x^4=0\] \[x(1-x^3)=0\] so x=1,0 but x cant be zero since 0 is not in the domain of the original function.

OpenStudy (zarkon):

and if x=1 then y=1

OpenStudy (anonymous):

so the only critical point is (1,1)?

OpenStudy (zarkon):

correct

OpenStudy (anonymous):

and its a saddle?

OpenStudy (zarkon):

looks like a min

OpenStudy (anonymous):

but i got that it was <0?

OpenStudy (zarkon):

what did you get for D?

OpenStudy (anonymous):

-1

OpenStudy (zarkon):

D=fxx(1,1)fyy(1,1)-[fxy(1,1)]^2

OpenStudy (anonymous):

ohh shoot i did it wrong

OpenStudy (zarkon):

I'm using my calculator and i get 3

OpenStudy (anonymous):

now i get 1

OpenStudy (anonymous):

fxx = 1, fyy=1 and fxy=0

OpenStudy (zarkon):

fxx=2/x^3 fyy=2/y^3 fxy=1

OpenStudy (anonymous):

ahhhh i didn't do chain rule, wow i'm on a role this morning

OpenStudy (anonymous):

okay got it thank you!

OpenStudy (zarkon):

yw